OpenStudy (anonymous):

find the vertical asymptotes of the rational function: f(x)= [x(x-1)]/(x^3+49x)... can someone give me the final answer??

OpenStudy (anonymous):

just find out when the denominator equals zero ex. x^3+49x=0, when ever you have something divided by zero you will get a vertical asymptote.

OpenStudy (anonymous):

yeah wat they said

OpenStudy (anonymous):

ok... x=0, +7i, -7i.. is it right??

OpenStudy (anonymous):

its 0=x^3+49x, so you solve for x, x doesnt equal zero

OpenStudy (anonymous):

wait nvm its weird to see i at an easier level. I suppose you are are good job!

OpenStudy (anonymous):

are right*

OpenStudy (anonymous):

yes:)... so the vertical is not equal x right??

OpenStudy (anonymous):

the vertical asymptote are those values of x, but the only real number is just zero. sorry I confused what you were doing. I thought you were setting x=0 then solving the denom.

OpenStudy (anonymous):

i were setting the denomial =0.. then i'm solving x.. i got x=0, 7i, -7i...

OpenStudy (anonymous):

yeap thats right haha good job linhphan

OpenStudy (anonymous):

thank you:)..